Circular No. 6424

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996aj IN ANONYMOUS GALAXY The Mount Stromlo Abell Cluster Supernova Search Team (cf. IAUC

6418) reports the discovery of an apparent supernova on CCD images taken by S. Chan on the Mount Stromlo Observatory 1.27-m telescope (+ Macho Camera). SN 1996aj is located at α = 13^{h}29^{m}06^{s}.82, δ = −29^{o}14′02′′.0, which is 4′′.45 west and 0′′.4 north of a galaxy situated near Abell 3559. A nearby star has position end figures 07^{s}.99, 20′′.9. SN 1996aj is visible on V and R CCD images taken on June 15 at mag ∼ 18.5 and was not detected on a poor frame obtained on May 28.

P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, write: “We obtained a spectrogram of SN 1996aj on June 19.3 UT with the Multiple Mirror Telescope. The spectrum shows that this is a type-Ia supernova between 10 and 20 days past maximum. The redshift of the supernova is between 0.10 and 0.11, making it background to the nearby Abell cluster 3559.”

SCORPIUS X-1 M. van der Klis and R. Wijnands, Astronomical Institute “Anton Pan-

nekoek”, University of Amsterdam; W. Chen, University of Maryland at College Park and Goddard Space Flight Center, NASA; F. K. Lamb and D. Psaltis, University of Illinois at Urbana-Champaign; E. Kuulkers, Astrophysics Division, European Space Agency and ESTEC; W. H. G. Lewin, Massachusetts Institute of Technology; B. Vaughan, California Institute of Technology; J. van Paradijs and S. Dieters, University of Alabama at Huntsville and University of Amsterdam; and K. Horne, University of St. Andrews, report: “Rossi X-ray Timing Explorer observations of Sco X-1 performed on May 25.0–25.2 UT show two kHz quasiperiodic oscillations (QPOs) similar to those observed in Feb. (IAUC 6319; van der Klis et al., Ap.J. Lett., submitted). Sco X-1 is at a lower accretion rate (as evidenced by the absence of 6-Hz QPO) than in the Feb. observations and the QPOs are at lower frequency (634 ± 2 and 926 ± 1 Hz). Their frequency difference is 292 ± 2 Hz, significantly larger than that in February (247 ± 3 Hz). This implies that the beat-frequency interpretation proposed for the two kHz QPOs in 4U 1728–34 (IAUC 6320, 6387; Strohmayer et al., Ap.J. Lett., submitted), where one QPO is due to the Keplerian rotation of the inner edge of the accretion disk and the other to the beat of this same Keplerian frequency with the neutron-star spin, can, for Sco X-1, be excluded.”

1996 June 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green